### v2

@@ Abstract
Choice overload – by which larger choice sets are detrimental to a chooser’s wellbeing – is potentially of great importance to the design of economic policy. Yet the current evidence on its prevalence is inconclusive. We argue that existing tests are likely to be underpowered and hence that choice overload may occur more often than the literature suggests. We propose more powerful tests based on richer data and characterization theorems for the Random Utility Model. These new approaches come with significant econometric challenges, which we show how to address. We apply our tests to new experimental data and find strong evidence of choice overload that would likely be missed using current approaches.
@@ 1 Introduction
The standard model of utility maximization tells us that increasing the set of available options can make the consumer no worse off, and may well improve their welfare. Since the pioneering study of Iyengar and Lepper (2000), [^1] See also Reibstein et al. (1975). a body of work in psychology has called this assumption into question. The umbrella term “choice overload” covers a number of phenomena by which larger choice sets seem to make people worse off. By now there are a huge number of studies investigating various aspects of choice overload; see Scheibehenne et al. (2010) and Chernev et al. (2015) for recent reviews and meta analyses.
Some of the output measures used to identify choice overload are hard to interpret using the classic tools of economic analysis; examples include a reduction in ex-post reported satisfaction or lower confidence that the right choice was made. Others fall very much in the realm of choice theory. In this paper we focus on the observation that larger choice sets may make people more likely to choose a “default” option.
Iyengar and Lepper (2000) provide an example of this type of behavior in their famous “jam” study. On different days a table was set up in a supermarket displaying a range of jams and offering a money-off coupon if a jam was bought. On “limited” choice days, six jams were available; on “extensive” choice days, customers faced 18 additional options. Strikingly, people were less likely to buy a jam on extensive choice days (2% vs 12% on limited choice days). [^2] These are the unconditional probabilities of buying jam. Typically the numbers quoted for this study are the probabilities of buying jam conditional on stopping at the table (3% vs 30%). However these numbers are more difficult to interpret as there may be selection into who stops at the table. Treating “do not buy” as the default choice option, this violates the Independence of Irrelevant Alternatives (IIA) axiom.
If choice overload could be well established it would have a number of important implications for economics. From a positive perspective, it would clearly violate the standard model of utility maximization. Moreover, it is not well explained by some of the more obvious behavioral theories. For example, most models of reference dependence assume that, for choices that share the same reference point, the decision maker will behave like a standard utility maximizer, ruling out choice overload type effects (Dean, 2008). Models of limited attention can also struggle to replicate the phenomenon, if they cannot explain why an overloaded individual does not simply ignore some of the available options. [^3] Though see Cattaneo et al. (2020, 2021) for examples of models of consideration sets, and de Lara and Dean (2024) for a model of sequential search that can accommodate choice overload-type behavior. This points to the need for models including more exotic elements such as regret (Sarver, 2008), rational contextual inference (Kamenica, 2008; Nocke and Rey, 2021) or decision avoidance (Beattie et al., 1994; Dean, 2008; Gerasimou, 2018). From a normative perspective, a key tenet of classic economics is that welfare is weakly improved by increasing the set of options available to the consumer. This assumption forms the basis of many policy recommendations, yet choice overload would call it into question.
@@
Unfortunately, current research into choice overload is inconclusive. Some direct replications of previous experiments have failed (Scheibehenne, 2008; Greifeneder et al., 2010). One recent meta-analysis concluded that the mean measured choice overload effect is zero (Scheibehenne et al., 2010). Another one (Chernev et al., 2015) concludes that whether or not choice overload exists may depend a lot on context.
The aim of this paper is to argue that existing studies are likely to underestimate the degree of choice overload of this type. [^4] Despite the above discussion, from now on we will use ‘choice overload’ to refer specifically to the phenomenon of choosing a default option more often in larger choice sets. Almost universally, a data set is said to exhibit choice overload only if the default is more likely to be chosen in a larger choice set than it is in some smaller choice set (see section 2). Yet intuitively, under the hypothesis of utility maximization one would expect the default to be chosen much less in larger choice sets, because the decision maker (DM) should be more likely to find something that they prefer to the default. This lack of power may potentially explain the inconsistent results from existing tests.
Two examples may further clarify this point.
@@
The following example is even more stylized but avoids distributional assumptions. Let the choice universe be $X=\{a,b,c,d\}$, where $d$ stands for default. Suppose $a,b,c$ are individually chosen with probability $20\%$ from choice sets $\{a,d\},\{b,d\},\{c,d\}$. Suppose also that the probability of active choice is $40\%$ for any of $\{a,b,d\},\{a,c,d\},\{b,c,d\}$. Assuming that indifference is not allowed, these probabilities can be reconciled with a Random Utility Model (RUM, precisely defined later); however, they imply that the preferences $a\succ d$, $b\succ d$, and $c\succ d$ occur in nonoverlapping subsets of the population. Thus, the active choice probability from $X$ must be at least $60\%$. If it is lower, we observe choice overload in the sense that observations cannot be explained in terms of a fully rational population, but they can be explained in terms of a population that is fully rational except that some individuals switch to default choice when encountering the grand choice set. This is true even if the active choice probability from $X$ strictly exceeds the active choice probability from any smaller set.
@@
In this paper, we describe new tests for choice overload. We consider a data set based on a finite set of alternatives $X$ that includes a default $d$. Observations of choice behavior are made from the grand set $X$, plus some subsets that always contain $d$. In each case we assume that the probability that $d$ is chosen is observed. We make use of two definitions of choice overload. The first is model free: we say that a data set exhibits choice overload if it violates monotonicity – i.e., the probability of choosing the default option increases as more alternatives are added to the choice set. The second is based on the null hypothesis of the random utility model (RUM) – i.e. there is a set of utility functions, a (choice set independent) probability distribution over that set, and the probability of choosing any option in any choice set is the probability of the set of utility functions that make that option maximal. As example 1.2 illustrates, consistency with RUM implies monotonicity, but not vise versa, so the latter definition provides a more sensitive test for choice overload. [^5] In general, it is well known that stochastic monotonicity is necessary but not sufficient for random utility. Example 1.2 clarifies that coarsening the domain of observations to active versus passive choice does not fundamentally change this. The advantage of the former is that it does not rely on the underlying assumption of random utility - monotonicity is also implied by a number of other models of stochastic choice, such as Manzini and Mariotti’s (2014) Random Consideration Set model.
If the data set consists only of choice probabilities from $X$ and a single subset $A$ (as is the case in Iyengar and Lepper (2000) and most other studies), then both definitions collapse to the conventional test: Assuming that only the probability of default choice is observed, the data set is consistent with monotonicity and RUM if and only if the probability of choosing $d$ from $X$ is weakly lower than from $A$. We therefore propose as a first modification of the standard approach that data should be collected from multiple subsets of $X$.
In this richer data, monotonicity implies that the choice of the default from $X$ is lower than the lowest probability of choosing $d$ across all observed choice sets $A$. We call this the Min bound. While easy to define, testing the Min bound is subtle because it is a test of multiple hypotheses. Even if monotonicity holds, if a finite number of observations are collected from many subsets there is a high probability that one will exhibit lower default choice than is observed from $X$. Essentially, the problem is the same as that of the “winners curse” – because the probability of default choice is observed with noise, the minimum of many observations is likely to be below the true minimum. To deal with this, we propose tests that take inspiration from closely related problems in the literature on moment inequalities.
Because monotonicity is necessary but not sufficient for RUM, our model-based definition of choice overload allows us to identify tighter bounds. We adapt the approach of McFadden and Richter (1991), who show that random choice data can be rationalized by the RUM if, and only if, a particular linear expression has a solution. We refer to the highest default choice probability in the large set for which such a solution exists as the RUM bound. The fact that it is choice frequencies, not underlying probabilities, that are observed presents a significant econometric challenge. We adapt the nonparametric test of RUM proposed by Kitamura and Stoye (2018), potentially including computational innovations by Smeulders et al. (2020) needed for applications with larger sets $X$.
@@
We apply our tests to a novel experimental data set based on choices from subsets of size $2$ and $3$ of $12$ different choice objects plus a default. The objects are verbally described sums of numbers as used in Caplin et al. (2011); see Figures 1 and 2 for a preview. In any given choice set, the subject could choose to stick with a default which provided \$3.50 for sure, or could switch to other available choices. Subjects were asked to chose from $10$ such sets, one of which was the default plus the entire set of $12$ other options –henceforth called the grand set– while the other $9$ were uniformly randomly selected from all possible sets with $1$ or $2$ options plus default.
A total of $2000$ subjects were recruited through Amazon’s Mechanical Turk platform. After removing subjects who failed a comprehension quiz, we were left with $1832$ subjects and $18320$ choices.
@@
Traditional approaches would be unlikely to find evidence of choice overload in our data set: the average default choice in small sets is higher than in the grand set, and only 15 of the 78 small sets have a default choice which is significantly lower than that of the grand set. Hence if, as in most work in the literature, data was collected from only a single or few small sets, one would be unlikely to detect overload. In contrast, even the most ‘brute force’ of our proposed tests –a potentially very conservative but finite-sample valid implementation of the Min bound– do detect it.
@@
Our test based on RUM provides an unexpected additional insight: While RUM is indeed rejected on the whole data set, it is also rejected if we remove data from choices in the grand set. This is because subjects are more likely to choose the default option in choice sets of size $3$ than their data from size $2$ sets and RUM would predict. Thus, our results indicate that choice overload type effects can start in choice sets which are much smaller than have previously been demonstrated.
@@ 2 Literature Review
Many studies have examined the phenomenon of choice overload from different perspectives, including different outcome measures and potential moderators. Two recent meta-analyses (Chernev et al., 2015; Scheibehenne et al., 2010) report results from 99 experiments in 53 papers and 63 experiments in 50 papers, respectively. Given these two reviews, our aim here is not to provide comprehensive coverage of the existing research. Instead, we make two points.
@@
First, to the best of our knowledge, no other paper has made use of the methodology we propose. Using the aforecited meta-analyses and Google Scholar, we identified 32 studies from 19 papers that use default choice as a measure of choice overload. Of these, 20 ask subjects to make choices from a single subset of the grand choice set. These studies can do no better than to compare the default choice probability in the small and large choice set. The remaining 12 studies ask subjects to make choices from multiple subsets of the large choice set, and so in principle could have implemented either the min bound or tests based on the RUM. However, all of them instead compare the average default choice across all small choice sets to that in the larger set, a measure which is necessary but sufficient for the aforementioned Min bound, which in turn is necessary but not sufficient for consistency with RUM. Applying our approach to the data from these previous studies (e.g., Chernev (2005), Gingras (2003), or White and Hoffrage (2009)) is an interesting avenue for future research.
@@
While not explicitly designed to test for choice overload, the experiment of Aguiar et al. (2023) does contain the type of data needed to perform our test. Choices were observed from all subsets of a grand set of six lotteries, with a default that is always available. The authors report finding no evidence of choice overload, and our tests confirm this result. [^6] While we performed the test on all their treatments, their ‘low’ complexity treatment is the best fit to apply the test. This is because, in the ‘medium’ and ‘high’ treatments, lottery prizes were expressed as sums, and different subjects saw different sums, arguably making them different choice alternatives. This may be due to the fact that the default alternative was chosen to be ‘obviously’ worse than the other available alternatives.
@@
The second point is that the veracity and scope of choice overload is far from established. Some direct replications have failed (Scheibehenne, 2008). One meta-analysis (Scheibehenne et al., 2010) finds a mean effect of set size on measures of choice overload to be zero, but noted a high degree of variance. A more recent analysis (Chernev et al., 2015) identifies four variables which can increase the incidence of choice overload: decision difficulty (for example due to time constraints), choice set complexity (for example due to hard to compare alternatives), preference uncertainty (for example because the decision maker is unsure how to aggregate their preferences across many dimensions), and decision goal (for example because the decision maker is not really committed to making a purchase). Chernev et al. (2015) argue that, taking these mediators into account, there is evidence of a robust choice overload effect.
We believe that the measures we introduce can contribute to this ongoing debate by providing more highly powered evidence on when choice overload is occurring. Low powered tests may in part explain some of the contradictory results in the existing literature – for example, some studies may by chance use a small choice set that leads to choice overload, while others do not. By making it easier to spot cases of overload, we anticipate movement towards a greater acceptance that choice overload does indeed occur.
@@ 3 Theory
We now present the theoretical underpinnings of our test for choice overload. We do so in three stages. First, we describe our tests under the assumption that we can perfectly observe the probabilities with which each alternative is chosen in each choice set. Second, we discuss the econometric techniques needed to handle the fact that our experimental data are sample frequencies, not population probabilities. Finally, we describe the steps necessary to implement these econometric tests, with additional considerations that proved unnecessary for our specific experiment relegated to Appendix B.
@@ 3.1 The Population-Level Testing Problem
Let $X$ be a finite set of alternatives and $d$ a default alternative contained in $X$. Let $\mathcal{D}\subset 2^{X}/\emptyset$ be a collection of choice sets, all of which contain the default alternative, and which includes the grand set $X$.
Our data set consists of observations of choice behavior in each set in $\mathcal{D}$ [^7] We follow established terminology in revealed preference theory by calling this a “data set,” but note that statistically speaking, these are population probabilities; inference for finite samples is considered in the next subsection. Specifically, we assume that we observe a function $p_{d}:\mathcal{D}\rightarrow[0,1]$, with the interpretation that $p_{d}(A)$ is the probability of choosing the default $d$ from choice set $A$.
Two notes on our data are in order. First, we assume that we observe the population probability with which the default is chosen in each choice set, but not that we can track individuals across choice problems; i.e., we observe a repeated cross section rather than a panel. This makes our approach applicable to many between-subject data sets that have this property. [^8] We note that the data we collect in our experiment is somewhat richer than this, as it contains multiple observations from the same individual. We maintain this assumption to increase the generalizability of our approach. Second, we assume that we observe only the probability with which the default was chosen in each choice set, not the probability with which specific non-default options are chosen. This is consistent with our desire to to focus on choice overload effects; the limited data only allows us to detect violations of utility maximization that occur to due too much or too little default choice from a set.
In order to design a test for choice overload, we first need to precisely define the term. Here we take two approaches. The first is to use an explicitly model-free definition, in the same way that risk aversion can be defined without reference to expected utility. An obvious candidate for such a definition is through violations of choice monotonicity, which states that the probability of choosing $d$ cannot increase as more options are added to the choice set.
@@
A data set $\{\mathcal{D},p\}$ satisfies monotonicity if, for any $A,B\in\mathcal{D}$ such that $A\subset B$,
@@
Specifically, one could declare choice overload to have occurred if there is a higher probability of default choice in the large choice set than in the smaller one. The canonical choice overload experiment, in which $\mathcal{D}=\{A,X\}$, applies this test.
More generally, one can define
@@
as the smallest observed probability of choosing $d$ in any set other than $X$. If $p_{d}(X)>p_{d}^{min}(X)$ then monotonicity is violated and the data set exhibits choice overload. We refer to $p_{d}^{min}(X)$ as the Min bound. Data that violates the Min bound is inconsistent with a number of models – most obviously RUM, but also the stochastic consideration set model of Manzini and Mariotti (2014) [^9] Other, more general models of stochastic consideration do allow for choice overload type effects - see for example Cattaneo et al. (2020, 2021) and models of reference dependent preferences such as Tversky and Kahneman (1991).
A second approach is to define choice overload as a violation of a specific model. Here, the most obvious candidate would be RUM, and we will work with it, although the basic idea would generalize to any model that we know how to test. Thus, call a data set stochastically rationalizable if it could have been generated by a RUM. Then we can think of it as revealing choice overload if it would be stochastically rationalizable except that the probability of default choice in larger choice sets is too large.
To illustrate this idea, consider the following definitions. Note that, while $\mathcal{D}$ in our application contains $X$, these definitions refer to a generic $\mathcal{D}$ that may or may not do so.
@@
A data set $\{\mathcal{D},p\}$ is consistent with RUM if there exist a finite collection $\mathcal{U}$ of one-to-one utility functions on $X$ and a probability distribution $\rho\in\Delta(\mathcal{U})$ such that, for every $A\in\mathcal{D}$,
@@
For any set $A\subseteq X$, we can then define a maximal bound on the choice of default using the default choice probabilities from the subsets of $A$ and consistency with RUM. The basic idea here goes back to Varian (1982, 1983): A counterfactual choice behavior is in the predictive bounds if, and only if, that choice behavior and previously observed ones are jointly rationalizable. Formally, we have:
@@
For a dataset $\{\mathcal{D},p\}$ and choice problem $A\subseteq X$, define
@@
defined on the choice sets $\mathcal{D}_{A}\equiv\{\tilde{A}\in\mathcal{D}:\tilde{A}\subset A\}$ is consistent with RUM.
@@
For most of this paper, we say that a data set exhibits choice overload if $p_{d}(X)>p^{RUM}_{d}(X)$. However, the definition of $p_{d}^{RUM}(\cdot)$ allows for choice overload to “kick in” for smaller choice sets and we will consider that later.
@@
Conceptually, the Min bound and the RUM bound deal with different reasons why observations from a single small choice set might not effectively identify choice overload. The first is heterogeneity in the quality of available items. Consider a grand choice set that consists of a number of not very appealing jam flavors (gooseberry, bark, salmon) and one extremely appealing flavor (strawberry). Absent any choice overload, we would expect to see high levels of default choice in small sets that do not include the strawberry jam, and low levels of default choice in both small sets that included the strawberry jam, and the grand set. Thus, if a researcher randomly selected for analysis a small choice set without the strawberry jam, it would make it very hard to spot choice overload. Collecting data on all small choice sets and applying the Min bound would help to solve this problem, because the researcher would observe choice from a small set in which strawberry jam was included.
A second issue is preference heterogeneity. Consider example 1.2 from the introduction. In this case, all alternatives are ex ante identical, so there is no heterogeneity in item quality. This in turn means that the default choice probability is the same in any choice set with the same cardinality. Thus, collecting (for example) data from multiple sets of size two and applying the Min bound does not help. However, as the example demonstrates, the RUM bound based on observations from choice sets of size two and three does put further constraints on the data. This is because the assumption that randomness is driven by preference heterogeneity puts extra structure on the allowed choice probabilities. Thus the RUM-based definition of choice overload gives potentially tighter bounds with which to identify choice overload. [^10] Note that, without preference heterogeneity, the RUM bound will equal the Min bounds. If all subjects have the same preferences, then in all small choice sets the default choice probability must be either zero or one. Thus, the Min bound must either be zero, if in some choice set the default is never chosen, or one, if the default is always chosen in every choice set. In both cases the RUM bound would be the same as the Min bound.

### v3

@@ Abstract
Choice overload – in which larger choice sets are detrimental to a chooser’s well-being – is potentially of great importance in the design of economic policy. Yet the current evidence on its prevalence is inconclusive. We argue that existing tests are likely to be underpowered and hence that choice overload may occur more often than the literature suggests. We propose more powerful tests based on richer data and characterization theorems for the Random Utility Model. These new approaches come with significant econometric challenges, which we show how to address. We apply our tests to new experimental data and find strong evidence of choice overload that would likely be missed using current approaches.
@@ 1 Introduction
Standard economic reasoning asserts that increasing the set of options cannot make the consumer worse off. Yet, starting with the pioneering study of Iyengar and Lepper (2000, see also ()), a growing body of work on choice overload – broadly speaking, the idea that people make worse choices in larger choice sets – has called this assumption into question. If true, choice overload would have important normative and positive implications for economics. It is inconsistent with utility maximization and many classic behavioral models. It also calls into question policy recommendations based on giving consumers as large a range of options as possible from which to choose.
@@
Unfortunately, current research on choice overload is inconclusive. Some direct replications of previous experiments failed (Scheibehenne, 2008; Greifeneder et al., 2010). One recent meta-analysis concludes that the mean measured choice overload effect is zero (Scheibehenne et al., 2010), and one (Chernev et al., 2015) concludes that its relevance may heavily depend on context. At the very least, there remains debate as to whether choice overload is a widespread and robust phenomenon.
In this paper, we argue that existing studies are likely to underestimate the extent of choice overload. A standard experimental paradigm for identifying overload is to compare the frequency of choosing a default option in small and large choice sets – for example, sticking with the default of not buying a jam in Iyengar and Lepper’s (2000, IL henceforth) classic study. [^1] While other measures of overload are used, such as ex post satisfaction and regret, Chernev et al. (2015) shows them to be highly correlated. A data set is said to exhibit choice overload if and only if the default is more likely to be chosen in a larger choice set than it is in the smaller choice set. Yet intuitively, under the hypothesis of utility maximization one would expect the default to be chosen much less in larger choice sets as the number of potentially attractive options grows. This lack of power may explain inconclusive results from existing tests.
Three examples may further clarify this point.
@@
Item heterogeneity Consider another variant of the IL example: Of the 24 jams IL use, 23 are variants of gooseberry jam, and one is strawberry jam. 10% of people like gooseberry jam, while everyone likes strawberry jam. This means that, absent choice overload, the probability of buying jam from the 24 jam set would be 100%. For a smaller choice set, it would be 100% if the strawberry jam were included or 10% otherwise. If the smaller set of jams selected by the researched happened not to include strawberry jam, then choice overload would have to be extremely strong to put active choice below the 10% benchmark.
Example 1.3.
@@
Tighter bounds Let the choice universe be $X=\{a,b,c,d\}$, where $d$ stands for default. Suppose $a,b,c$ are individually chosen across a population with probability $20\%$ from choice sets $\{a,d\},\{b,d\},\{c,d\}$. Suppose also that the probability of non-default choice is $40\%$ from any of $\{a,b,d\},\{a,c,d\},\{b,c,d\}$. Assuming no indifferences, these probabilities can be reconciled with a Random Utility Model (RUM, precisely defined later); however, they imply that the preferences $a\succ d$, $b\succ d$, and $c\succ d$ occur in disjoint subsets of the population. Thus, non-default choice probability from $X$ must be at least $60\%$, even though the largest such probability observed in smaller sets is “only” $40\%$.
@@
We introduce new tests for choice overload for a data set in which default versus nondefault choice is observed for a grand set $X$ and numerous subsets thereof (always including the same default option). We define choice overload as a situation in which the default is chosen too often in $X$ relative to some benchmark. We identify two tests based on two different benchmarks. The first equates choice overload with violations of monotonicity – that is, the probability of choosing the default option increases as choice sets expand. This is the approach previously used in the literature, though the structure of our data still allows for an improved test. The second is based on presupposing a RUM – that is, choices can be explained by a (choice set independent) probability distribution over a set of utility functions. As example 1.3 illustrates, consistency with RUM implies monotonicity, but not vise versa, so the latter definition provides a more sensitive test for choice overload. [^2] In general, it is well known that stochastic monotonicity is necessary but not sufficient for random utility. Example 1.3 clarifies that coarsening the domain of observations to active versus passive choice does not change this.
If the data consists only of choice probabilities from $X$ and a single subset $A$, then both definitions coincide and reduce to the current best practice. We therefore propose that data should be collected from multiple subsets of $X$. In this richer data, monotonicity implies that choice probability of the default from $X$ is less frequent than from any other choice set $A$. We call this the Min bound. While easy to define, testing this restriction is subtle because it is a test of multiple hypotheses; we propose an approach that takes inspiration from closely related problems in the literature on moment inequalities.
We next observe that our model-based definition of choice overload implies tighter bounds. We lean on classic work by McFadden and Richter (1991) to characterize the highest default choice probability in the large set that is consistent with the data from subsets and RUM, which we call the RUM bound. To implement this condition in practice, we adapt Kitamura and Stoye’s (2018, see also Smeulders et al. (2021)) nonparametric test of RUM.
@@
We apply all tests to a novel experimental data set of choices from subsets of size $2$ and $3$ of $12$ different choice objects plus a default. The objects are verbally described sums of numbers as used in Caplin et al. (2011); see Figure LABEL:fig:1 for a preview. Subjects were asked to chose from $10$ such sets, one of which was the default plus the entire set of $12$ other options (henceforth called the grand set).
@@
Traditional approaches would be unlikely to find evidence of choice overload in our data set: The average default choice in small sets is higher than in the grand set, and only 15 of the 78 small sets have a default choice which is significantly lower than that of the grand set. In contrast, even the most brute force of our proposed tests – a potentially very conservative but finite-sample valid implementation of the Min bound – do detect it.
@@
Our test based on RUM provides an unexpected additional insight: While RUM is indeed rejected on the whole data set due to choice overload, it is also rejected if we remove data from choices in the grand set. This is because subjects are more likely to choose the default option in choice sets of size $3$ than RUM would predict given their data from size $2$ sets. Thus, our results indicate that choice overload type effects can start in choice sets which are much smaller than was previously demonstrated.
We conclude this section with two important clarifications. First, the more sensitive RUM-based test comes at the expense of additional assumptions – for example, monotonicity is also implied by Tversky and Kahneman’s (1991) model of reference dependent preferences. So our refined test is really a test of choice overload for a population that is supposed to otherwise adhere to a specific model of rationality. We believe, however, that this is in line with what is typically meant by choice overload. Secondly, it is important to emphasize that our notion of choice overload is “thin” or behavioral. That is, we think of choice overload as a specific feature of choice behavior: that the default is chosen too often in larger sets. We do not commit to a theory of what is going on “under the hood.” For example, the increase in default choice set in larger sets observed in our data is plausibly explained by a rational inattention heuristic. If true, we would describe this a a case of rational inattention causing choice overload, rather than an alternative description of what is going on. We naturally think that “deep” theories of the mechanisms generating choice overload are important and we hope that our work aids future researchers in uncovering them; however, our main aim in this paper is not to substantively advance that inquiry but to provide tools with which to identify choice overload.
@@ 2 Literature Review
Two recent meta-analyses (Chernev et al., 2015; Scheibehenne et al., 2010) report results from 99 experiments in 53 papers and 63 experiments in 50 papers, respectively. We refer to these reviews and are content to just make two points.
@@
First, very few current studies have generated the data needed to perform our tests. Using the aforementioned meta-analyses and Google Scholar, we identified 32 studies from 19 papers that use default choice as a measure of choice overload. Of these, 20 ask subjects to make choices from a single subset of the grand choice set. These studies can do no better than to compare the default choice probability in the small and large choice set. The remaining 12 studies ask subjects to make choices from multiple subsets of the large choice set, but only collect a small number of choices from each. As a result, they instead compare the average default choice across all small choice sets to that in the larger set, a measure which is necessary but not sufficient for the aforementioned Min bound, which in turn is necessary but not sufficient for consistency with RUM. While applying our approach to data from these previous studies is an interesting avenue for future research, the small number of observations in each small set make our tests hard to implement. [^3] We note that our tests can be applied to any data – not just experimental – where a large sample of default vs non-default choice is observed from a choice set and multiple subsets. For example, data on health insurance plan choices, such as that used in Abaluck and Gruber (2023), often has variation in the specific plans offered as well as the number offered which could be used for our analysis. We leave such applications of our tests for future research.
@@
Although not explicitly designed to test for choice overload, the experiments of Aguiar et al. (2023) contain the type of data needed to perform our test. Choices were observed from all subsets of a grand set of six lotteries, with a default that is always available. The authors report no evidence of choice overload, and our tests confirm this result. This may be due to the fact that the default alternative in their experiment was chosen to be obviously worse than the other available alternatives.
@@
The second point is that the veracity and scope of choice overload is far from established. Some direct replications have failed (Scheibehenne, 2008). One meta-analysis (Scheibehenne et al., 2010) finds the mean effect of set size on measures of choice overload to be zero, but notes a high variance. A more recent analysis (Chernev et al., 2015) identifies four variables which can increase the incidence of choice overload: decision difficulty (for example due to time constraints), choice set complexity (for example due to hard-to-compare alternatives), preference uncertainty (for example because the decision maker is unsure how to aggregate their preferences across many dimensions), and decision goal (for example because the decision maker is not really committed to making a purchase).
Finally, our testing problem is related to that studied in Kono et al. (2023). [^4] A recent paper by Turansick (2025) is more distantly related. He considers characterizations of RUM when not all choice problems are observable, but for those that are all probabilities are observed. We both test RUM when not all choice probabilities from a choice set are observed. While also taking inspiration from McFadden and Richter (1991), Kono et al. identify a different set of conditions using the Block-Marschak polynomials. This approach is elegant and may have computational benefits, but there are some reasons why our approach is more directly applicable to the task at hand. First, while Kitamura and Stoye (2018) (and we) use RUMs as motivation, they (and we) really just test whether vectorized choice probabilities lie in a certain polyhedral cone. In application to RUM, the vertices of this cone are defined by “choice types” that are conventional utility maximizers, but one can analogously use this approach to test the “random utility extension” of other models of behavior (we do this with generalizations of RUM that capture choice overloaded behavior). In addition, Kono et al. (2023) assume that the collection of observable choice sets is closed under set expansion; we do not need this and it does not hold in our data.
@@ 3 Theory
We now present the theoretical underpinnings of our test. We initially assume that we can perfectly observe default choice probabilities for each choice set; in a second step, we develop the econometric tools required because our actual data are finite samples. Additional computational considerations that proved unnecessary for our experiment are relegated to supplementary materials.
@@ 3.1 The Population-Level Testing Problem
Let $X$ be a finite set of alternatives and $d$ a default alternative contained in $X$. Let $\mathcal{D}\subset 2^{X}/\emptyset$ be a collection of choice sets, all of which contain $d$; in most of what follows, and in our experimental data, $\mathcal{D}$ will contain $X$.
Suppose initially that we observe a function $p_{d}:\mathcal{D}\rightarrow[0,1]$, where $p_{d}(A)$ is the probability of choosing the default $d$ from choice set $A$. We assume that we observe the population probability with which the default is chosen in each choice set, but not that we can track individuals across choice problems. This makes our approach applicable to many between-subject data sets. Crucially, we also assume that we observe only the probability with which the default was chosen in each choice set, not the probability with which specific non-default options are chosen. This is consistent with our desire to to focus on choice overload effects; the limited data only allows us to detect violations of utility maximization that occur to due too much or too little default choice from a set, and the tests we will define will identify the former violations as ‘choice overload’.
We next consider two possible definitions of choice overload. The first one equates it with violations of choice monotonicity.
@@
Probabilities $p_{d}$ satisfy monotonicity if, for any $A,B\in\mathcal{D}$ such that $A\subset B$,
@@
The canonical choice overload experiment, in which $\mathcal{D}=\{A,X\}$, tests this condition. More generally, one can define
@@
as the smallest observed probability of choosing $d$ in any set other than $X$. Monotonicity is violated iff $p_{d}(X)>p_{d}^{min}(X)$. We therefore refer to $p_{d}^{min}(X)$ as the Min bound. We will say that data that violates this condition as exhibiting choice overload with respect to the Min bound. Such data is inconsistent with a number of models – most obviously RUM, but also models of reference dependent preferences such as Tversky and Kahneman (1991). [^5] This is also true for the stochastic consideration set model of Manzini and Mariotti (2014), a special case of RUM. Other more general models of stochastic consideration allow for choice overload type effects – see, for example, Cattaneo et al. (2020, 2021).
A second approach is to define choice overload as a violation of a specific model. Here, the most obvious candidate is RUM, and we will work with it, although the basic idea would generalize to any model that we know how to test. Thus, call a data set stochastically rationalizable if it could have been generated by a RUM. Then we can think of it as revealing choice overload if it would be stochastically rationalizable except that the probability of default choice in larger choice sets is too large.
The basic idea is that one would declare a data set to exhibit choice overload if the default is chosen too often in the grand set $X$, given the choice probabilities from the smaller sets and the constraints imposed by RUM. One could implement this either by testing for the validity of RUM both including or excluding data from $X$, or by computing counterfactual bounds on $p_{d}(X)$ as the set of all default choice probabilities on $X$ that would be rationalizable jointly with the other observed probabilities.
To illustrate this idea, consider the following definitions.
@@
Probabilities $p_{d}$ are consistent with RUM if there exist a finite collection $\mathcal{U}$ of one-to-one utility functions on $X$ and a probability distribution $\rho\in\Delta(\mathcal{U})$ such that, for every $A\in\mathcal{D}$,
@@
For any $A\subseteq X$, we can then define a maximal bound on the choice of default using the default choice probabilities from the subsets of $A$ and consistency with RUM. The basic idea here goes back to Varian (1982, 1983): A counterfactual choice behavior is in the predictive bounds if, and only if, that choice behavior and previously observed ones (in our case, behavior on small choice sets) are jointly rationalizable. Formally:
@@
Define
@@
on choice sets $\mathcal{D}_{A}\equiv\{\tilde{A}\in\mathcal{D}:\tilde{A}\subseteq A\}$ being jointly consistent with RUM.
@@
For most of this paper, we say that a data set exhibits choice overload according to the RUM bound if $p_{d}(X)>p^{RUM}_{d}(X)$. However, the definition of $p_{d}^{RUM}(\cdot)$ allows for choice overload to “kick in” for smaller choice sets and we will consider that later.
@@
The Min and RUM bounds speak to different reasons why observations from a single small choice set might not effectively identify choice overload. The first is heterogeneity in the quality of available items, as illustrated in example 1.2 . Consider a grand choice set that consists of a number of not very appealing jam flavors and one extremely appealing flavor (strawberry, say). Absent any choice overload, we would expect high levels of default choice in small sets that do not include the strawberry jam, and low levels of default choice in both small sets that included the strawberry jam, and the grand set. Thus, if a researcher randomly selected for analysis a small choice set without the strawberry jam, it would make it very hard to spot choice overload. Collecting data on all small choice sets and applying the Min bound would address this problem.
The second issue is preference heterogeneity. Consider Example 1.3: Here, all alternatives are equally likely to be attractive. Default choice probability is the same in all choice sets of the same size, and so a simple application of the Min bound will not increase power to detect choice overload. However, the RUM bound based on observations from choice sets of size two and three further constrains choice probabilities. (In a perfectly homogeneous population, the bounds coincide.)